# Emergence of a High-Dimensional Abstraction Phase in Language Transformers

### v2

@@ 1 Introduction
Taken together, our experiments suggest that all analyzed transformer architectures develop, in the intermediate layers, a high-dimensional representation encoding complex and abstract linguistic information. The results of this processing are stored in representations which are then used, possibly through a process of incremental refinement, to predict the next token. Besides providing new insights on the inner workings of Transformer-based LMs, these results have implications for tasks such as layer-based fine-tuning, model pruning and model stitching.
@@ 2 Related work
The notion that nominally complex, high-dimensional objects can be described using few degrees of freedom is not new: it underlies, for instance, a number of popular dimensionality reduction methods such as standard PCA (Jolliffe, 1986). But, while PCA is linear, the data manifold need not be: as such, general nonlinear methods have been proposed to estimate the topological dimension, or manifold dimension, of point clouds (see Campadelli et al. (2015) for a survey). As neural representations tend to constitute nonlinear manifolds across modalities (Ansuini et al., 2019; Cai et al., 2021), we use a state-of-the-art nonlinear ID estimation method, the Generalized Ratios Intrinsic Dimension Estimator (GRIDE) (Denti et al., 2022), which we further describe in Section 3.4.
@@
Complementary to parameter ID, a number of works focus on the ID of representations in LM activation space. In particular, Cai et al. (2021) were the first to identify low-dimensional manifolds in the contextual embedding space of (masked) LMs. Balestriero et al. (2023) linked representational ID to the scope of attention and showed how toxicity attacks can exploit this relationship. Tulchinskii et al. (2023) demonstrated that representational ID can be used to differentiate human- and AI-generated texts. Cheng et al. (2023) established a relation between representational ID and information-theoretic compression, also showing that dataset-specific ID correlates with ease of fine-tuning.
Closer to our aims, Valeriani et al. (2023) studied the evolution of ID across layers for vision and protein transformers, with a preliminary analysis of a single language transformer tested on a single dataset. Like us, they found that ID develops in different phases, with a consistent early peak followed by a valley and less consistent markers of a second peak. We greatly extend their analysis of linguistic transformers by investigating the functional role of the main ID peak in five distinct LMs and a number of new analyses and experiments. Importantly, our converging evidence suggests that, in language transformers, semantic processing of the input first takes place under the first ID peak, contra the preliminary evidence by Valeriani and colleagues that this crucial phase is the low-ID “valley” between the peaks.
@@ 3.1 Models
OPT, Pythia, and OLMo make public their pre-training datasets, which are a combination of web-scraped text, code, online forums such as Reddit, books, research papers, and encyclopedic text (see Zhang et al. (2022); Gao et al. (2020); Soldaini et al. (2024), respectively, for details). The pre-training datasets of Llama-3 and Mistral are likely similar, though they remain undisclosed at the time of submission.
@@ 3.2 Data
Since we focus on model behavior in-distribution, we compute observables using three corpora that proxy models’ pre-training data (all accessed through HuggingFace): Bookcorpus (Zhu et al., 2015); the Pile (Gao et al. (2020); precisely, the 10k-document subsample available on HuggingFace) and WikiText-103 (Merity et al., 2017). From each corpus, we randomly extract 5 non-overlapping partitions of 10k 20-token sequences (sequence length is counted according to the number of typographic tokens in the text of each corpus, as made available on HuggingFace). [^2] We replicated the Pile-based Pythia and OPT experiments with sequences extended to 128 tokens. We obtained very similar results, confirming that the sequences we are using cover the typical contextual spans encoded in model representations. We do not constrain the sequences to have any particular structure: in particular, their final element is not required to coincide with the end of a sentence. We additionally generated 5 partitions of 10k 20-token sequences from shuffled versions of the same corpora. The latter respect the source corpus unigram frequency distribution, but syntactic structure and semantic coherence are destroyed.
@@ 3.3 Probing and downstream tasks
If there is a relation between ID and abstract linguistic processing, then ID may also predict ease-of-transfer to downstream NLP tasks. To test this claim, we consider two such tasks in a zero-shot setting, sentiment classification of film reviews (Maas et al., 2011) and toxicity classification (Adams et al., 2017), for which we train binary linear classifiers on each hidden layer representation (training details in Appendix J). Note that we do not fine-tune the layers themselves, as this might change their IDs, which is the very property we want to relate to task performance.
@@ 3.4 Intrinsic Dimension
In almost every real-world system, the ID depends on the scale, i.e., the size of neighbourhood at which the data is analyzed. In particular, at small scales, the true dimensionality of the manifold is typically hidden by that of data noise. At very large scales, the ID estimate can also be erroneous, due, for instance, to the curvature of the manifold (Facco et al., 2017; Denti et al., 2022). For this reason, in order to obtain a reliable and meaningful ID estimation, a proper scale analysis is necessary. This is typically performed by varying the amount of the neighbours considered in estimating the ID and looking for an interval of scales in which the estimate is approximately stable (see Appendix C for further details). To this aim, we opted for the generalized ratios intrinsic dimension estimator (GRIDE) of Denti et al. (2022), which extends the commonly used [^3] E.g., Ansuini et al. (2019); Tulchinskii et al. (2023); Valeriani et al. (2023). TwoNN estimator of Facco et al. (2017) to general scales.
In GRIDE, the fundamental ingredients are ratios $\mu_{i,2k,k}=r_{i,2k}/r_{i,k}$, where $r_{i,j}$ is the Euclidean distance between point $i$ and its $j$-th nearest neighbour. Under local uniform density assumptions, the $\mu_{i,2k,k}$ follow a generalised Pareto distribution $f_{\mu_{i,2k,k}}(\mu)=\frac{d(\mu^{d}-1)^{k-1}}{B(k,k)\mu^{d(2k-1)+1}}$, where $B(\cdot,\cdot)$ is the beta function. By assuming the empirical ratios $\mu_{i,2k,k}$ to be independent for different points, one obtains the ID by numerically maximizing the mentioned likelihood.
@@ 3.5 Quantifying the relative information content of different representations
In comparing the layers’ representation spaces, we also considered three alternative measures: Doimo et al. (2020)’s neighborhood overlap measure, also used by Valeriani et al. (2023), Representational Similarity Analysis (Kriegeskorte et al., 2008), and linear CKA (Kornblith et al., 2019). Only the last measure resulted in interpretable results, broadly coherent with those obtained with $\Delta$ (Appendix H).
@@ 4.1 Emergence of a central high-dimensionality phase
Figure 2 reports $\Delta(l_{i}\to l_{first})$ and $\Delta(l_{i}\to l_{last})$ for Llama, OPT and Pythia (with the remaining models in Appendix E). We observe that the ID peak largely overlaps with a peak in $\Delta(l_{i}\to l_{first})$. Here, a large value of $\Delta$ implies that sequences which are nearest neighbours at the ID-peak are distant from each other in the input layer. That is, the ID peak layers no longer contain the information encoded in the initial representation of the sequence, suggesting they instead capture higher-level information. Meanwhile, we note that the $\Delta(l_{i}\to l_{last})$ profiles are not clearly related to the first ID peak, and we leave their analysis to future work.
@@ 4.2 Language processing during the high-dimensionality phase
Figure 5: Linguistic knowledge probing performance $\pm$ 2 SDs across 5 random seeds is shown with the ID for Llama, OPT, and Pythia (left to right). Row (a) corresponds to surface-form tasks Sentence Length and Word Content, where probe performance decreases through the ID peak. Row (b) corresponds to syntactic and semantic tasks Bigram Shift, Coordination Inversion and Odd Man Out, where probe performance for all tasks attains maximum (or close) within the ID peak. This suggests the ID peak marks abstract, and not surface, representations of the input.
@@
Figure 5(b) reports accuracy for Llama, Pythia, and OPT (results for the remaining models in Appendix I) on the syntactic and semantic probe tasks. As described in detail in Appendix I, these tasks require a model to detect whether a sentence is syntactically well-formed (Bigram Shift), whether it is semantically coherent (Odd Man Out) and whether it describes a pair of events or states in a plausible order (Coordination Inversion). Despite variation across tasks and models, we observe that, in general, asymptotic accuracy is reached during the ID peak phase. Again, this implies that, for OPT, the asymptote is reached later. For OLMo (Appendix I), we observe in some cases a late accuracy peak, related to the second ID expansion phase that this model undergoes. In general, however, once top accuracy is reached, performance stays quite constant across layers, suggesting that the expressivity of high dimensionality permits rich linguistic representaiton of the inputs, but, once this information is extracted, it is propagated across subsequent layers. This is intuitive, as high-level linguistic information is useful to the network for its ultimate task of next-token prediction. Appendix D confirms the same patterns for a smaller and a larger Pythia model.
@@ 5 Conclusion
Our paper focused on detecting broad qualitative patterns. As it is the first study of its kind, we wanted to establish some basic high-level results. Without the expectations set by our high-level investigation, we cannot make precise predictions to be tested in more specialized mechanistic-interpretability studies. We note moreover that, while high-level, our conclusions are supported by very consistent converging evidence from a variety of models, corpora and experiments (even OLMo, that is to some extent the “outlier” model, displays the ID peak and the relevant associated properties). Future work should establish a clearer causal connection between intrinsic dimension and language processing, via layer ablations studies and by constraining the dimensionality of different layers during language processing.
@@
From a more applied perspective, we see various ways in which our discovery of a consistent high-ID phase could be useful. ID profiles emerge as significant blueprints of model behaviour that could be used as proxies of model quality. ID information can be used for model pruning, or to choose which layers to fine-tune, or for model stitching and other model-interfacing operations. These are all exciting directions for future work.
@@ 6 Reproducibility statement
We uploaded corpora, code and a readme document to reproduce our experiments as supplementary materials, and we will make them publicly available upon acceptance. We report our compute usage in Appendix A, and describe all publicly available resources we used in Appendix B. Appendices Appendix C, Appendix I and Appendix J specify relevant hyperparameter choices for ID computation, probing task and transfer task implementation, respectively.
@@ C.1 Scale analysis
In this section we explicitly show how to perform a scale analysis in order to select a proper and meaningful scale when computing the ID. In fact, as explained in the main text, the ID estimation might be affected by undesirable effects that are always present when dealing with real-world datasets and that can hide the true dimensionality of the manifold underlying the data. Such effects include the presence of noise, which typically affects small scales, and density variations and manifold curvature, which, instead, tend to be observed at a larger scale (Denti et al., 2022). It is thus recommended to choose an intermediate scale, where the ID estimate might be stable across different neighbourhood sizes, and it is less likely to be affected by the aforementioned spurious factors. For this reason, it is necessary to see how the ID changes when varying the neighbourhood size taken into account when performing the ID computation. We rely on the GRIDE estimator, which allows to explicitly select the number of neighbours considered. In particular, the rank of the first nearest neighbour used to compute the distance ratio is a hyperparameter, and we refer to the chosen value for this hyperparameter as the scale in what follows.
We performed scale analysis for each model and corpus. An example is provided in Figure C.1, which shows the results for Pythia on Pile. We plot the ID estimate for increasing scales (i.e., number of neighbours, x-axis in Figure C.1) for some representative layers. The rank of the neighbour considered is explored by means of powers of 2 in order to look at regions where the variation of the ID would be noticeable. The true ID is likely to lie at a scale in which the ID estimate approximately plateaus (Denti et al., 2022), which is marked by the highlighted region. For simplicity, per model-corpus combination, we choose one scale for all layers: in this particular example, we choose $k=2^{5}$. In general, the optimal neighbourhood size tends to be around $k=32$ for all (model, corpus) combinations, allowing us to reliably compare them (see Table C.1 for all $k$). Once the scale is chosen for each (model, corpus) combination, we plot the scale-adjusted ID estimates (see, e.g., Figure C.3).
@@ C.2 Additional results
Figure C.3 displays the evolution of estimated ID, scale-adjusted, per layer for all corpora and models. While the magnitude of ID differs across corpora, with ID on Bookcorpus lower than those of the Pile and Wikitext for all models, all corpora exhibit the characteristic high-ID peak, and at nearly the same onset. The dampened Bookcorpus peak IDs, which are still significantly above the corresponding shuffled ID peaks, might be explained by the fact that this corpus is entirely made of novels, and it is thus the less in-domain of the corpora we explored.
@@ Appendix E Information imbalance with respect to first/last layer
Figure E.1 shows averaged $\Delta(l_{i}\to l_{first})$ (gray) and $\Delta(l_{i}\to l_{last})$ for Mistral and OLMo. Recall that the closer $\Delta(A\to B)$ is to 0, the more predictive $A$’s local neighborhood structure is of $B$’s. As expected, $\Delta(l_{i}\to l_{first})$ generally increases with $i$ as we go deeper in the layers; the reverse is true for $\Delta(l_{i}\to l_{last})$. However, $\Delta(l_{i}\to l_{first})$ appears to locally peak and plateau around the representational ID peak; that is, the ID expansion marks a phase of low predictivity from the intermediate layer to the input. OLMo’s pattern is not as clear, and we might observe a second local information imbalance peak in proximity to the second ID peak characterizing this model.
Figure E.2 shows, in the spirit of the Information Plane (Shwartz-Ziv & Tishby, 2017), the $\Delta(l_{i}\to l_{last})$ plotted against $\Delta(l_{first}\to l_{i})$ for all models across the layers. The trajectory reveals the dynamic evolution of input and output informativity along the residual stream.
@@
Figure E.2: Information Plane: $\Delta(l_{i}\to l_{last})$ vs. $\Delta(l_{first}\to l_{i})$. For each LM, the information imbalance $\Delta$ from the $i$th to the last layer is plotted against $\Delta$ from the first to the $i$th layer, where lower $\Delta$ means more informative. Each point on the curve corresponds to a single layer $i$; for all models, the first layer (gold star) begins on the y-axis, $\Delta(l_{first}\to l_{i})=0$, and the trajectory through the layers ends on the x-axis, $\Delta(l_{i}\to l_{last})=0$. Consistently across models, ID peak layers (red) mark a changepoint in processing patterns. Pre-peak layers show a rapid increase along the x-axis but slow decrease along the y-axis, indicating a rapid departure from the inputs while not learning much about the outputs. This reflects the phase of semantic abstraction. Post-peak layers’ information dynamics vary by model, though all are marked by a final rapid decrease in $\Delta(l_{i}\to l_{last})$ and a decrease in $\Delta(l_{first}\to l_{i})$. This suggests a return to surface-level information processing in order to predict the next token.

### v3

@@ 1 Introduction
Taken together, our experiments suggest that all analyzed transformer architectures develop, in the intermediate layers, a high-dimensional representation encoding abstract linguistic information. The results of this processing are stored in representations which are then used, possibly through a process of incremental refinement, to predict the next token. Besides providing new insights on the inner workings of Transformer-based LMs, these results have implications for tasks such as layer-based fine-tuning, model pruning and model stitching.
@@ 2 Related work
The notion that nominally complex, high-dimensional objects can be described using few degrees of freedom is not new: it underlies, for instance, a number of popular dimensionality reduction methods such as standard (linear) PCA (Jolliffe, 1986). But, while PCA is linear, the data manifold need not be: as such, general nonlinear methods have been proposed to estimate the topological dimension, or manifold dimension, of point clouds (see Campadelli et al. (2015) for a survey). As neural representations tend to constitute nonlinear manifolds across modalities (Ansuini et al., 2019; Cai et al., 2021), we use a state-of-the-art nonlinear ID estimation method, the Generalized Ratios Intrinsic Dimension Estimator (GRIDE) (Denti et al., 2022), which we further describe in Section 3.4.
@@
Complementary to parameter ID, a number of works focus on the ID of representations in LM activation space. In particular, Cai et al. (2021) were the first to identify low-dimensional manifolds in the contextual embedding space of (masked) LMs. Balestriero et al. (2023) linked representational ID to the scope of attention and showed how toxicity attacks can exploit this relationship. Tulchinskii et al. (2023) demonstrated that representational ID can be used to differentiate human- and AI-generated texts. Cheng et al. (2023) established a relation between representational ID and information-theoretic compression, also showing that dataset-specific ID correlates with ease of fine-tuning. Yin et al. (2024) showed that the local ID of specific inputs in specific layers can be used to diagnose model truthfulness.
Closer to our aims, Valeriani et al. (2023) studied the evolution of ID across layers for vision and protein transformers, with a preliminary analysis of a single language transformer tested on a single dataset (a similarly preliminary analysis is also provided by (Yin et al., 2024)). Like us, Valeriani and colleagues found that ID develops in different phases, with a consistent early peak followed by a valley and less consistent markers of a second peak. We greatly extend their analysis of linguistic transformers by investigating the functional role of the main ID peak in five distinct LMs. Importantly, our converging evidence suggests that, in language transformers, semantic processing of the input first takes place under the early ID peak, contra the preliminary evidence by Valeriani and colleagues that this crucial phase is the low-ID “valley” between the peaks.
@@ 3.1 Models
OPT, Pythia, and OLMo make public their pre-training datasets, which are a combination of web-scraped text, code, online forums such as Reddit, books, research papers, and encyclopedic text (see Zhang et al. (2022); Gao et al. (2020); Soldaini et al. (2024), respectively, for details). The pre-training datasets of Llama-3 and Mistral are likely similar, though they remain undisclosed at the time of publication.
@@ 3.2 Data
Since we focus on model behavior in-distribution, we compute observables using three corpora that proxy models’ pre-training data (all accessed through HuggingFace): Bookcorpus (Zhu et al., 2015); the Pile (Gao et al. (2020); precisely, the 10k-document subsample available on HuggingFace) and WikiText-103 (Merity et al., 2017). From each corpus, we sample, without replacement, a total of 50k distinct 20-token sequences (sequence length is counted according to the number of typographic tokens in the text of each corpus). [^2] We replicated the Pile-based Pythia and OPT experiments with sequences extended to 128 tokens. We obtained very similar results, confirming that the sequences we are using cover the typical contextual spans encoded in model representations. We then divide these samples into partitions of 10k sequences each, which we use for the experiments. We do not constrain the sequences to have any particular structure: in particular, their final element is not required to coincide with the end of a sentence. We additionally generate 5 partitions of 10k 20-token sequences from shuffled versions of the same corpora. That is, we first randomize the order of the tokens in each corpus, and then proceed as with the non-randomized versions. The shuffled samples respect the source corpus unigram frequency distribution, but syntactic structure and semantic coherence are destroyed.
@@ 3.3 Probing and downstream tasks
If there is a relation between ID and abstract linguistic processing, then ID may also predict ease-of-transfer to downstream NLP tasks. To test this claim, we consider two such tasks, sentiment classification of film reviews (Maas et al., 2011) and toxicity classification (Adams et al., 2017), for which we train binary linear classifiers on each hidden layer representation (training details in Appendix J). Note that we do not fine-tune the layers themselves, as this might change their IDs, which is the very property we want to relate to task performance.
@@ 3.4 Intrinsic Dimension
In almost every real-world system, the ID depends on the scale, i.e., the size of neighbourhood at which the data is analyzed. In particular, at small scales, the true dimensionality of the manifold is typically hidden by that of data noise. At very large scales, the ID estimate can also be erroneous, due, for instance, to the curvature of the manifold (Facco et al., 2017; Denti et al., 2022). For this reason, in order to obtain a reliable and meaningful ID estimation, a proper scale analysis is necessary. This is typically performed by varying the amount of the neighbours considered in estimating the ID and looking for an interval of scales in which the estimate is approximately stable (see Appendix C for further details). To this aim, we opted for the generalized ratios intrinsic dimension estimator (GRIDE) of Denti et al. (2022), which extends the commonly used [^3] E.g., Ansuini et al. (2019); Tulchinskii et al. (2023); Valeriani et al. (2023). TwoNN estimator of Facco et al. (2017) to general scales. We selected GRIDE because it allows probing, in a rigorous framework, the dependence of ID on scale. While the original TwoNN estimator assumes local uniformity up to the 2nd nearest neighbor, GRIDE relaxes this assumption to produce unbiased ID estimates up to the $2k$th nearest neighbor ($k$ being the scale).
In GRIDE, the fundamental ingredients are ratios $\mu_{i,2k,k}=r_{i,2k}/r_{i,k}$, where $r_{i,j}$ is the Euclidean distance between point $i$ and its $j$-th nearest neighbour. Under local uniform density assumptions, the $\mu_{i,2k,k}$ follow a generalised Pareto distribution $f_{\mu_{i,2k,k}}(\mu)=\frac{d(\mu^{d}-1)^{k-1}}{B(k,k)\mu^{d(2k-1)+1}}$ that depends on the intrinsic dimension $d$ of the manifold, where $B(\cdot,\cdot)$ is the beta function. By assuming the empirical ratios $\mu_{i,2k,k}$ to be independent for different points, one obtains an estimate of the intrinsic dimension $\hat{d}$ by numerically maximizing the mentioned likelihood.
@@ 3.5 Quantifying the relative information content of different representations
In comparing the layers’ representation spaces, we also considered three alternative measures broadly used in deep net representation analysis (see Sucholutsky et al., 2024; Williams, 2024, for recent surveys): Doimo et al. (2020)’s neighborhood overlap measure, also used by Valeriani et al. (2023), Representational Similarity Analysis (Kriegeskorte et al., 2008), and linear CKA (Kornblith et al., 2019). Only the last measure resulted in interpretable results, broadly coherent with those obtained with $\Delta$ (Appendix H).
@@ 4.1 Emergence of a central high-dimensionality phase
Figure 2 reports $\Delta(l_{i}\to l_{first})$ and $\Delta(l_{i}\to l_{last})$ for Llama, OPT and Pythia (with the remaining models in Appendix E, that also presents further analysis). We observe that the ID peak largely overlaps with a peak in $\Delta(l_{i}\to l_{first})$. Here, a large value of $\Delta$ implies that sequences which are nearest neighbours at the ID-peak are distant from each other in the input layer. That is, the ID peak layers no longer contain the information encoded in the initial representation of the sequence, suggesting they instead capture higher-level information. Meanwhile, we note that the $\Delta(l_{i}\to l_{last})$ profiles are not clearly related to the first ID peak, and we leave their analysis to future work.
@@ 4.2 Language processing during the high-dimensionality phase
Figure 5: Linguistic knowledge probing performance $\pm$ 2 SDs across 5 random seeds is shown with the ID profile for Llama, OPT, and Pythia (left to right). Row (a) corresponds to surface-form tasks Sentence Length and Word Content, where probe performance decreases through the ID peak. Row (b) corresponds to syntactic and semantic tasks Bigram Shift, Coordination Inversion and Odd Man Out, where probe performance for all tasks attains maximum (or close) within the ID peak. This suggests the ID peak marks abstract, and not surface, representations of the input.
@@
Figure 5(b) reports accuracy for Llama, Pythia, and OPT (results for the remaining models in Appendix I) on the syntactic and semantic probe tasks. As described in detail in Appendix I, these tasks require a model to detect whether a sentence is syntactically well-formed (Bigram Shift), whether it is semantically coherent (Odd Man Out) and whether it describes a pair of events or states in a plausible order (Coordination Inversion). Despite variation across tasks and models, we observe that, in general, asymptotic accuracy is reached during the ID peak phase. Again, this implies that, for OPT, the asymptote is reached later. For OLMo (Appendix I), we observe in some cases a late accuracy peak, related to the second ID expansion phase that this model undergoes. In general, however, once top accuracy is reached, performance stays quite constant across layers, suggesting that the expressivity of high dimensionality permits rich linguistic representation of the inputs, but, once this information is extracted, it is propagated across subsequent layers. This is intuitive, as high-level linguistic information is useful to the network for its ultimate task of next-token prediction. Appendix D confirms the same patterns for a smaller and a larger Pythia model.
@@ 5 Conclusion
Our paper focused on detecting broad qualitative patterns. We found very consistent converging evidence from a variety of models, corpora and experiments (even OLMo, that is to some extent the “outlier” model, displays the ID peak and the relevant associated properties). Future work should establish a clearer causal connection between intrinsic dimension and language processing, via layer ablations studies and by constraining the dimensionality of different layers during language processing.
@@
From a more applied perspective, we see various ways in which our discovery of a consistent high-ID phase could be useful. ID profiles emerge as significant blueprints of model behaviour that could be used as proxies of model quality. ID information can be used for model pruning, or to choose which layers to fine-tune, or for model stitching and other model-interfacing operations, such as training LM-based encoding models of the brain (Antonello & Cheng, 2024). These are all exciting directions for future work.
@@ 6 Reproducibility statement
Please find corpora, code and a readme document to reproduce our experiments at https://github.com/chengemily1/id-llm-abstraction . We report our compute usage in Appendix A, and describe all publicly available resources we used in Appendix B. Appendix C, Appendix I and Appendix J specify relevant hyperparameter choices for ID computation, probing task and transfer task implementation, respectively.
7 Acknowledgments
We thank the members of the COLT group, Mor Geva, the ICLR reviewers, and the area chair for useful feedback. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 101019291). This paper reflects the authors’ view only, and the funding agency is not responsible for any use that may be made of the information it contains. Additionally, D.D. received support by the project “Supporto alla diagnosi di malattie rare tramite l’intelligenza artificiale”(CUP: F53C22001770002).
@@ C.1 Scale analysis
In this section we explicitly show how to perform a scale analysis in order to select a proper and meaningful scale when computing the ID. As explained in the main text, ID estimation might be affected by undesirable effects that are always present when dealing with real-world datasets and that can hide the true dimensionality of the manifold underlying the data. Such effects include the presence of noise, which typically affects small scales, and density variations and manifold curvature, which, instead, tend to be observed at a larger scale (Denti et al., 2022). It is thus recommended to choose an intermediate scale, where the ID estimate might be stable across different neighbourhood sizes, and it is less likely to be affected by the aforementioned spurious factors. For this reason, it is necessary to see how the ID changes when varying the neighbourhood size taken into account when performing the ID computation. We rely on the GRIDE estimator, which allows to explicitly select the number of neighbours considered. In particular, the rank of the first nearest neighbour used to compute the distance ratio is a hyperparameter, and we refer to the chosen value for this hyperparameter as the scale in what follows.
We performed scale analysis for each model and corpus. An example is provided in Figure C.1, which shows the results for Pythia on Pile. We plot the ID estimate for increasing scales (i.e., number of neighbours, x-axis in Figure C.1) for some representative layers. The rank of the neighbour considered is explored by means of powers of 2 in order to look at regions where the variation in ID is noticeable. The true ID is likely to lie at a scale in which the ID estimate approximately plateaus (Denti et al., 2022), which is marked by the highlighted region. For simplicity, per model-corpus combination, we choose one scale for all layers: in this particular example, we choose $k=2^{5}$. In general, the optimal neighbourhood size tends to be around $k=32$ for all (model, corpus) combinations, allowing us to reliably compare them (see Table C.1 for all $k$). Once the scale is chosen for each (model, corpus) combination, we plot the scale-adjusted ID estimates (see, e.g., Figure C.3).
@@ C.2 Additional results
Figure C.3 displays the evolution of estimated ID, scale-adjusted, per layer for all corpora and models. While the magnitude of ID differs across corpora, with ID on Bookcorpus lower than those of the Pile and Wikitext for all models, all corpora exhibit the characteristic high-ID peak, and at nearly the same onset. The dampened Bookcorpus peak IDs, which are still significantly above the corresponding shuffled ID peaks, might be explained by the fact that this corpus is entirely made of novels (and lower-cased), and it is thus the less in-domain of the corpora we explored.
@@ Appendix E Information imbalance with respect to first/last layer
Figure E.1 shows averaged $\Delta(l_{i}\to l_{first})$ (gray) and $\Delta(l_{i}\to l_{last})$ (brown) for Mistral and OLMo. Recall that the closer $\Delta(A\to B)$ is to 0, the more predictive $A$’s local neighborhood structure is of $B$’s. As expected, $\Delta(l_{i}\to l_{first})$ generally increases with $i$ as we go deeper in the layers; the reverse is true for $\Delta(l_{i}\to l_{last})$. However, $\Delta(l_{i}\to l_{first})$ appears to locally peak and plateau around the representational ID peak; that is, the ID expansion marks a phase of low predictivity from the intermediate layer to the input. OLMo’s pattern is not as clear, and we might observe a second local information imbalance peak in proximity to the second ID peak characterizing this model.
Figure E.2 shows, in the spirit of the Information Plane (Shwartz-Ziv & Tishby, 2017), the $\Delta(l_{i}\to l_{last})$ plotted against $\Delta(l_{first}\to l_{i})$ for all models across the layers. The trajectory reveals the dynamic evolution of input and output informativity along the residual stream. Consistently across models, ID peak layers (red) mark a change point in processing patterns. Pre-peak layers show a rapid increase along the x-axis but slow decrease along the y-axis, indicating a rapid departure from the inputs while not learning much about the outputs. This reflects the phase of semantic abstraction. Post-peak layers’ information dynamics vary by model, though all are marked by a final rapid decrease in $\Delta(l_{i}\to l_{last})$ and a decrease in $\Delta(l_{first}\to l_{i})$. This suggests a return to surface-level information processing in order to predict the next token.
@@
Figure E.2: $\Delta(l_{i}\to l_{last})$ vs. $\Delta(l_{first}\to l_{i})$. For each LM, the information imbalance $\Delta$ from the $i$th to the last layer is plotted against $\Delta$ from the first to the $i$th layer, where lower $\Delta$ means more informative. Each point on the curve corresponds to a single layer $i$; for all models, the first layer (gold star) begins on the y-axis, $\Delta(l_{first}\to l_{i})=0$, and the trajectory through the layers ends on the x-axis, $\Delta(l_{i}\to l_{last})=0$. Consistently across models, ID peak layers (red) mark a changepoint in processing patterns. Pre-peak layers show a rapid increase along the x-axis but slow decrease along the y-axis, indicating a rapid departure from the inputs while not learning much about the outputs. This reflects the phase of semantic abstraction. Post-peak layers’ information dynamics vary by model, though all are marked by a final rapid decrease in $\Delta(l_{i}\to l_{last})$ and a decrease in $\Delta(l_{first}\to l_{i})$. This suggests a return to surface-level information processing in order to predict the next token.